OpenStudy (anonymous):

What is the relationship between exponential functions and logarithmic functions? Then describe how their graphs are similar and different.

OpenStudy (anonymous):

well, the exponential function is actually defined as the inverse of the logarithmic function. they are the same graph just reflected over the line y = x (it's just the inverse)

OpenStudy (dumbcow):

they are inverse functions e^(lnx) = x ln(e^x) = x